OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. I got y - 2 = -3(x + 7)??

jimthompson5910 (jim_thompson5910):

close, but you have your signs mixed up

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) y - y1 = -3(x - x1) y - (-2) = -3(x - 7) y + 2 = -3(x - 7)

OpenStudy (anonymous):

ok i see what you did thanks alot

jimthompson5910 (jim_thompson5910):

yw